Can you help me solve (Given: n is odd )(Prove: n^2 is odd): Can you help me solve (Given: n is odd )(Prove: n^2 is odd)
Answer by stanbon(19670)

Can you help me solve (Given: n is odd )(Prove: n^2 is odd)
-------------
If n is odd, n = 2x+1 where x is an integer.
-------
Then n^2 = (2x+1)^2 = 4x^2+4x + 1 = 2(2x^2+2x) +1
Since x is an integer, 2x^2+2x is an integer.
So, 2(2x^2+2x)+1 follows the form 2(integer)+1
So (2x+1)^2 is odd.
Therefore n^2 is odd.
Cheers,
Stan H.